OpenStudy (anonymous):

Write an expression that represents the area of a rectangle with length 6a^2b and width 3ab. (For any rectangle with l and width w, A=lw)

OpenStudy (anonymous):

so wats the problem??? you just have to put values. and get the answer for area

OpenStudy (anonymous):

The problem is I'm not sure how to do it. It's difficult for me.

OpenStudy (anonymous):

The area is equal to the length times the width. We have the length and the width, so we multiply them together: \[6a^2 b \times 3ab\] Since we have a 6 and 3, we can multiply them together to get 18. We multiply a^2 by a, following the rule that multiplying the same base means you add the exponents. So a^2 times a = a^(2+1) = a^3. The same rule with the b gives you b^2. So our answer, when we put it all together, is: \[18 a^3 b^2\]

OpenStudy (anonymous):

Wow that's actually really simple. Thanks for the help mate. :) So let's say the problem was like 5a^2 x 4a^2 the same rules apply? like would the answer be 20a^4?

OpenStudy (anonymous):

Yep, exactly. When you multiply two things with the same base, you just add the exponents. If you divide, you subtract the exponents.

OpenStudy (anonymous):

Thanks. Cheers! :)

OpenStudy (anonymous):

You're welcome, glad I could help.